# A Volumetric Privacy Measure for Dynamical Systems With Bounded Disturbance

### v4

@@ 3.3 Properties of the Interval Inference Attack
The inference attack exhibits certain properties. For instance, the radius of the uncertainty $\mathcal{Y}_{k|k}$, i.e., $p_{k|k}^{y}=\overline{Y}_{k|k}-\underline{Y}_{k|k}$ is bounded by a function of the radius of the disturbance and the observation set, as stated below.
@@
For any $k\geqslant 1$, the radius of $\mathcal{Y}_{k|k}$ satisfies
$p_{k|k}^{y}\leqslant$ $\left(\left|A_{3}\right|+\left|A_{4}\right|\left|A_{2}^{-1}\right|+\left|A_{4}\right|\left|A_{2}^{-1}A_{1}\right|\right)\overline{p}^{x}$
$+\left|A_{4}\right|\left|A_{2}^{-1}B_{1}\right|p^{w,x}_{k}+\left|B_{2}\right|p^{w,y}_{k},$ (41)
where $\overline{p}^{x}\geqslant p^{m,x}_{j|j}$ for any $j\geqslant 0$, $p^{m,x}_{k|k}$, $p^{w,x}_{k}$ and $p^{w,y}_{k}$ are radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$, respectively, and $|A|$ is a matrix where each element is the absolute value of the corresponding element in $A$, i.e., $|A|=\left[|a_{i,j}|\right]$.
Proof 3.4.
@@
Since the volume of $\mathcal{Y}_{k|k}$ is the sum of $p_{k|k}^{y}$, $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ is also bounded by a function of the radius of the observation set $\mathcal{M}_{k|k}^{x}$. Consequently, if $\mathcal{M}_{k|k}^{x}$ is small, then the privacy level would be low, and the adversary retains little uncertainty after performing the inference attack.
Furthermore, by comparing the predicted and posterior uncertainty sets, as given by (5) and (10), the amount of uncertainty reduction can be quantified by $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$, where
$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}\setminus\mathcal{Y}_{k|k},$
is the difference between $\mathcal{Y}_{k|k-1}$ and $\mathcal{Y}_{k|k}$. Since $\mathcal{Y}_{k|k}$ is a subset of $\mathcal{Y}_{k|k-1}$, $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ can be computed with
$\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)\!.$ (42)
Therefore, to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$, it is equivalent to reduce the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ since the amount of prior uncertainty $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$ is fixed at time $k$. As shown in the next theorem, the amount of uncertainty reduction, i.e., $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$, is bounded by the new information extracted from $\mathcal{M}^{x}_{k|k}$.
@@
$\!\!\!\!\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\!=\!\left\|\Psi\!\left(A_{3}\right)\!\Delta\mathcal{X}_{k-1|k}\oplus\Psi\!\left(A_{4}\right)\!\Delta\mathcal{Y}_{k-1|k}\right\|_{1}\!,\!\!$ (43)
satisfying
$\mathrm{Vol}$ $\left(\Delta\mathcal{Y}_{k|k}\right)\geqslant 2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1},$
$\!\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)\!\!\leqslant$ $\!\left\|\!A_{3}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!+\!\left\|\!A_{4}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)\!,$
where $\Delta\mathcal{X}_{k-1|k}\!\!=\!\!\mathcal{X}_{k-1|k-1}\!\setminus\!\mathcal{X}_{k-1|k}$, $\Delta\!\mathcal{Y}_{k-1|k}\!\!=\!\!\mathcal{Y}_{k-1|k-1}\!\setminus\!\mathcal{Y}_{k-1|k}$ are the adversary’s uncertainty reduction of $X_{k-1}$ and $Y_{k-1}$, and $\left\|c_{k|k}^{y}\!-\!c_{k|k-1}^{y}\right\|_{1}$ quantifies the difference in central estimation with and without considering $\mathcal{X}_{k|k}$.
Proof 3.5.
@@
According to Theorem 9, the reduction of uncertainty $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is highly correlated with the amount of information that the adversary extracts from the observation set $\mathcal{M}_{k|k}^{x}$. Furthermore, with (42) and Theorem 9, we can show the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k-1|k}\right)$ is bounded in the following lemma.
@@
The privacy level can be bounded with the following inequalities,
$\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k-1}\!\right)\!-\!\left\|\!A_{3}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!-\!\left\|\!A_{4}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)$
$\leqslant\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k}\!\right)\!\leqslant\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$
As a result, we can reduce the extracted information $\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)$ and $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right)$ to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ via designing proper observation set $\mathcal{M}^{x}_{k|k}$. Moreover, when the privacy level is high, the adversary’s ability to update its central estimate $c_{k|k}^{y}$ is also limited, as indicated by the small value of $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$. Based on this observation, $\mathcal{M}^{x}_{k|k}$ can also be designed to hinder accurate central estimate updates, further improving the privacy level.
@@ 4.1 The Structure of Privacy Filter
which contains all possible public states that can be reached from any states in $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$. Therefore, $\mathcal{X}_{k|k-1}$ is the maximum observation set $\mathcal{M}^{x}_{k|k}$ that the filter can release, i.e., $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$. To maintain the high data utility, the observation set has to satisfy the following constraint,
$\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right)\leqslant\epsilon^{x}.$
@@
Note that $\mathcal{S}^{x}_{k|k}$ is randomly generated as a subset of $\mathcal{X}_{k|k-1}$. In practice, it can be sufficiently small; for instance, it may contain only the true public state $x_{k}$. Consequently, recovering $\mathcal{S}^{x}_{k|k}$ could lead to potential privacy leakage and should therefore be avoided. In the following, we show that an attacker cannot recover $\mathcal{S}^{x}_{k|k}$ by inverting the optimization problem $\mathbf{P_{1}}$ due to the randomization operation.
@@
As a result, the proposed privacy filter exhibits the following properties: (1) In the absence of inference attacks, the filter output $\mathcal{M}^{x}_{k|k}$ satisfies the utility constraint. (2) In the presence of the inference attack described in (6)–(12), the filter output maximizes the privacy level. (3) The filter is robust against reverse attacks that attempt to recover the sufficiently small set $\mathcal{S}_{k|k}^{x}$.
Moreover, since the computational complexity of both the inference attack in (6)–(12) and the volume computation increases with the complexity of the set representations, a trade-off exists between the achievable privacy enhancement of the proposed filter and its computational cost. While more sophisticated set representations could improve the accuracy of privacy evaluation, for efficiency and clarity, we next present a concrete design based on the interval approximation described in Sec. 3.2.
@@ 4.2 Randomization
$\mathrm{Vol}\left(\mathcal{S}^{x}_{k|k}\right)=$ ${\beta}_{k}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}+{\alpha}_{k}\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}$
@@ 4.3 Privacy Filter Optimization
The privacy filter optimization problem $\mathbf{P_{1}}$ is equivalent to the following linear programming
@@
Consequently, we can solve the linear programming problem $\mathbf{P_{2}}$ to obtain the optimal observation set that defends the system against the inference attack defined in Section 2.3.
@@ 5 Numerical Verification
We first plot the trajectories of system states and their interval tubes in Fig.3 under the optimal privacy filter design for different values of $\epsilon^{x}$. We also use the central point of the posterior intervals as one of the possible testing estimation, e.g., $\frac{\overline{X}_{k|k}+\underline{X}_{k|k}}{2}$ for $\mathcal{X}_{k|k}$. The pink areas in these figures represent the uncertainty sets of system states. As shown in Fig.3, when $\epsilon^{x}=0.01$ is small, the adversary’s uncertainty about the private production rate is small, and its central estimation closely matches the actual production rate. However, when $\epsilon^{x}$ increases to $0.5$, the utility of the inventory information decreases slightly, but this leads to higher uncertainty of the inference attack, causing the adversary’s central estimation of the production rate to become less accurate. This observation numerically verifies Theorem 9, confirming that a higher privacy level prevents the adversary from refining its incorrect central estimation. Therefore, the proposed privacy filter effectively reduces the privacy leakage of the production rate, though at the cost of introducing some inaccuracies in the inventory information.
@@
The privacy-utility trade-off was evaluated by plotting the average privacy level of the production rate against the average utility of the inventory over 100 random trajectories with horizon 100, as illustrated in Fig. 5. For a fair and clear comparison, the privacy level and utility values for the truncated Gaussian mechanism are normalized to the range $[0,1]$, and its same scaling parameters are applied to the other two mechanisms. The results demonstrate that an increase in inventory utility corresponds to a reduction in the privacy level of the production rate, thereby confirming the intrinsic trade-off between data utility and privacy protection. Besides, the quantization mechanism and the truncated Gaussian mechanism have similar performance in privacy-utility trade-off. Note that these two mechanisms assume static domains of states, while the sets of states in systems subject to disturbance can be estimated and described more precise as addressed by the inference attack. The proposed volumetric approach estimates the time-varying private state set and adjusts the output accordingly.
Furthermore, as discussed in [31], if the volume (e.g., interval length) of domain satisfies certain conditions, the truncated Gaussian mechanism ensures differential privacy. However, given the volume constraint, the shape of the public state set can be arbitrary, and certain shapes may lead to substantial volumetric leakage of the private state after inference attacks. The proposed volumetric method explicitly accounts for this by considering the geometry of the set based on the assumed inference attack, not only its volume. Therefore, it achieves higher privacy levels while maintaining lower data distortion compared with the two static mechanisms.

### v5

@@ 3.4 Properties of the Interval Inference Attack
The inference attack exhibits several key properties. In particular, the radius of the uncertainty set $\mathcal{Y}_{k|k}$, $p_{k|k}^{y}=\overline{Y}_{k|k}-\underline{Y}_{k|k}$, is bounded by a function of the radius of the disturbance and the observation set, as formalized below.
@@
For any $k\geq 1$, the radius of $\mathcal{Y}_{k|k}$ satisfies
$p_{k|k}^{y}$ $\leq\Big(|A_{3}|+|A_{4}||A_{2}^{-1}|+|A_{4}||A_{2}^{-1}A_{1}|\Big)\overline{p}^{x}$
$\quad+|A_{4}||A_{2}^{-1}B_{1}|p_{k}^{w,x}+|B_{2}|p_{k}^{w,y},$ (45)
where $\overline{p}^{x}\geq p^{m,x}_{j|j}$ for any $j\geq 0$, $p^{m,x}_{k|k}$, $p^{w,x}_{k}$, and $p^{w,y}_{k}$ are the radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{W}_{k}^{x}$, and $\mathcal{W}_{k}^{y}$, respectively, and $|A|$ denotes the matrix with elementwise absolute values, i.e., $|A|=[|a_{i,j}|]$.
Proof 3.5.
@@
Since the volume of $\mathcal{Y}_{k|k}$ is given by the product of $2p_{k|k}^{y}$ over all dimensions, $\mathrm{Vol}(\mathcal{Y}_{k|k})$ is similarly bounded by a function of the radius of the observation set $\mathcal{M}_{k|k}^{x}$. Consequently, a small $\mathcal{M}_{k|k}^{x}$ implies low privacy, and the adversary retains limited uncertainty after performing the inference attack.
Furthermore, by comparing the predicted and posterior uncertainty sets, as in (13) and (18), the reduction of uncertainty can be quantified using the surrogate measure $\overline{\mathrm{Vol}}(\Delta\mathcal{Y}_{k|k})$, where
$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}\setminus\mathcal{Y}_{k|k}.$ (46)
Since $\mathcal{Y}_{k|k}\subseteq\mathcal{Y}_{k|k-1}$, the surrogate volume can be computed as
$\overline{\mathrm{Vol}}(\Delta\mathcal{Y}_{k|k})=\overline{\mathrm{Vol}}(\mathcal{Y}_{k|k-1})-\overline{\mathrm{Vol}}(\mathcal{Y}_{k|k}).$ (47)
Thus, increasing the privacy level $\overline{\mathrm{Vol}}(\mathcal{Y}_{k|k})$ is equivalent to reducing the amount of uncertainty reduction $\overline{\mathrm{Vol}}(\Delta\mathcal{Y}_{k|k})$, since the prior uncertainty $\overline{\mathrm{Vol}}(\mathcal{Y}_{k|k-1})$ is fixed at time $k$. As shown in the next theorem, the amount of uncertainty reduction, is bounded by the new information extracted from $\mathcal{M}^{x}_{k|k}$.
@@
$\!\!\!\!\overline{\mathrm{Vol}}\left(\Delta\mathcal{Y}_{k|k}\right)\!=\!\left\|\Psi\!\left(A_{3}\right)\!\Delta\mathcal{X}_{k-1|k}\oplus\Psi\!\left(A_{4}\right)\!\Delta\mathcal{Y}_{k-1|k}\right\|_{1}\!,\!\!$ (48)
with bounds
$\overline{\mathrm{Vol}}$ $\left(\Delta\mathcal{Y}_{k|k}\right)\geqslant 2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1},$
$\!\overline{\mathrm{Vol}}\!\left(\Delta\mathcal{Y}_{k|k}\right)\!\!\leqslant$ $\!\left\|\!A_{3}\!\right\|_{1}\!\overline{\mathrm{Vol}}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!+\!\left\|\!A_{4}\!\right\|_{1}\!\overline{\mathrm{Vol}}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)\!,$
where $\Delta\mathcal{X}_{k-1|k}\!\!=\!\!\mathcal{X}_{k-1|k-1}\!\setminus\!\mathcal{X}_{k-1|k}$, $\Delta\!\mathcal{Y}_{k-1|k}\!\!=\!\!\mathcal{Y}_{k-1|k-1}\!\setminus\!\mathcal{Y}_{k-1|k}$, and $\|c_{k|k}^{y}-c_{k|k-1}^{y}\|_{1}$ quantifies the change in the central estimate due to the observation $\mathcal{X}_{k|k}$.
Proof 3.6.
@@
Combining (47) and Theorem 9, the privacy level $\overline{\mathrm{Vol}}(\mathcal{Y}_{k|k})$ can be bounded as follows.
@@
The privacy level satisfies
$\overline{\mathrm{Vol}}\!\left(\!\mathcal{Y}_{k|k-1}\!\right)\!-\!\left\|\!A_{3}\!\right\|_{1}\!\overline{\mathrm{Vol}}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!-\!\left\|\!A_{4}\!\right\|_{1}\!\overline{\mathrm{Vol}}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)$
$\leqslant\overline{\mathrm{Vol}}\!\left(\!\mathcal{Y}_{k|k}\!\right)\!\leqslant\overline{\mathrm{Vol}}\left(\mathcal{Y}_{k|k-1}\right)-2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$
As a result, we can reduce the extracted information $\overline{\mathrm{Vol}}\left(\Delta\mathcal{X}_{k-1|k}\right)$ and $\overline{\mathrm{Vol}}\left(\Delta\mathcal{Y}_{k-1|k}\right)$ to increase the privacy level $\overline{\mathrm{Vol}}\left(\mathcal{Y}_{k|k}\right)$ via designing proper observation set $\mathcal{M}^{x}_{k|k}$. Moreover, if the privacy level is high, the adversary’s ability to update its central estimate $c_{k|k}^{y}$ is also limited, as indicated by the small value of $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$. Based on this observation, $\mathcal{M}^{x}_{k|k}$ can also be designed to hinder accurate central estimate updates, further improving the privacy level.
@@ 4.1 The Structure of Privacy Filter
which contains all possible public states that can be reached from any states in $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$. Therefore, $\mathcal{X}_{k|k-1}$ is the maximum observation set $\mathcal{M}^{x}_{k|k}$ that the filter can release, i.e., $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$. To maintain high data utility, the uncertainty set $\mathcal{X}_{k|k}$ must satisfy the following constraint:
$\mathrm{Vol}\left(\mathcal{X}_{k|k}\right)\leq\epsilon^{x},$
where $\epsilon^{x}>0$ specifies the desired upper bound on the uncertainty volume. Since $\mathcal{X}_{k|k}$ is a subset of the observation set $\mathcal{M}_{k|k}^{x}$, this constraint can be equivalently enforced on the larger set, $\mathrm{Vol}\left(\mathcal{M}_{k|k}^{x}\right)\leq\epsilon^{x}$, which simplifies the design of the observation set while ensuring that the utility requirement is satisfied.
@@
Note that $\mathcal{S}^{x}_{k|k}$ is randomly generated as a subset of $\mathcal{X}_{k|k-1}$ and can be made sufficiently small in practice. For instance, the random set $\mathcal{S}^{x}_{k|k}$ may contain only the true public state $x_{k}$. According to Lemma 8, a sufficiently small observation set could lead to potential privacy leakage. Therefore, recovering $\mathcal{S}^{x}_{k|k}$ from $\mathbf{P_{1}}$ must be avoided. In the following, we demonstrate that an attacker cannot recover $\mathcal{S}^{x}_{k|k}$ by inverting the optimization problem $\mathbf{P_{1}}$, owing to the randomization mechanism embedded in the privacy filter.
@@
As a result, the proposed privacy filter exhibits the following properties: (1) In the absence of inference attacks, the filter output $\mathcal{M}^{x}_{k|k}$ satisfies the utility constraint. (2) In the presence of the inference attack described in (14)–(20), the filter output maximizes the privacy level. (3) The filter is robust against reverse attacks that attempt to recover the sufficiently small set $\mathcal{S}_{k|k}^{x}$, thereby reducing the risk of privacy compromise from adversaries exploiting structural vulnerabilities.
Moreover, the computational complexity of both the inference attack in (14)–(20) and the volume computation increases with the complexity of the set representations. Consequently, a trade-off exists between the achievable privacy enhancement of the proposed filter and its computational cost. While more sophisticated set representations may improve the accuracy of privacy evaluation, for efficiency and clarity, we next present a concrete design based on the interval approximation described in Sec. 3.3.
In addition, as discussed in Sec.3.3, the surrogate measure $\overline{\mathrm{Vol}}(\cdot)$ bounds the volume of interval, and could simplify the computation complexity. We adopt the surrogate volumetric measure $\overline{\mathrm{Vol}}(\cdot)$ in the optimization to further reduce computational complexity, and we verify that using this surrogate measure still leads to improved privacy levels in Sec. 5.
@@ 4.2 Randomization
$\overline{\mathrm{Vol}}\left(\mathcal{S}^{x}_{k|k}\right)=$ ${\beta}_{k}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}+{\alpha}_{k}\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}$
@@ 4.3 Privacy Filter Optimization
The privacy filter optimization problem $\mathbf{P_{1}}$ with the surrogate privacy measure $\overline{\mathrm{Vol}}\left(\mathcal{Y}_{k|k}\right)$ and utility measure $\overline{\mathrm{Vol}}\left(\mathcal{M}^{x}_{k|k}\right)$ is equivalent to the following linear programming
@@
Consequently, we can solve the linear programming problem $\mathbf{P_{2}}$ to obtain the optimal observation set that defends the system against the inference attack defined in Section 3.
@@ 5 Numerical Verification
We first plot the trajectories of the system states and their corresponding interval tubes in Fig. 3 under the optimal privacy filter design for different values of $\epsilon^{x}$. The shaded pink areas represent the interval tubes, i.e., the uncertainty sets of the system states, which quantify the range of values that an adversary can infer. As shown in Fig. 3, for $\epsilon^{x}=0.01$, the adversary’s uncertainty about the private production rate is small. However, increasing $\epsilon^{x}$ to $0.5$ slightly reduces the utility of the inventory information but significantly enlarges the adversary’s uncertainty.
To further illustrate the effectiveness of the privacy filter, we consider one possible adversary estimate based on the central points of the posterior intervals. For the public state $\mathcal{X}_{k|k}$, the adversary’s central estimate is given by $\frac{\overline{X}_{k|k}+\underline{X}_{k|k}}{2}$. As shown in Fig. 3, when $\epsilon^{x}$ increases from $0.01$ to $0.5$, the adversary’s central estimate of the production rate becomes significantly less accurate, whereas the central estimate of the inventory information remains accurate. This observation numerically confirms Theorem 9, showing that a higher privacy level prevents the adversary from refining an incorrect central estimate. Hence, the proposed privacy filter effectively mitigates the leakage of production rate information, while introducing a controlled loss of accuracy in the inventory information.
@@
We evaluate the privacy-utility trade-off by plotting the average privacy level of the production rate against the average utility of the inventory in Fig. 5. For a fair and clear comparison, the privacy level and utility values for the truncated Gaussian mechanism are normalized to the range $[0,1]$, and the same scaling parameters are applied to the other two mechanisms. The results demonstrate that an increase in inventory utility corresponds to a reduction in the privacy level of the production rate, thereby confirming the intrinsic trade-off between data utility and privacy protection.
As discussed in [31], if the interval length, i.e., volume, of domain satisfies certain conditions, the truncated Gaussian mechanism ensures differential privacy. However, given the volume constraint, the shape of the public state set can be arbitrary, and certain shapes may lead to substantial volumetric leakage of the private state after inference attacks. The proposed volumetric method explicitly accounts for this by considering the geometry of the set based on the assumed inference attack, not only its volume. Therefore, it achieves higher privacy levels while maintaining lower data distortion compared with the two static mechanisms.